# The efficiency of black hole formation via collisions in stellar systems:

### v1

@@ 4 Summary and conclusions
In this paper, we have investigated how the efficiency of black hole formation via collisions in stellar systems depends on the ratio of stellar over critical mass. The importance of the collision timescale in the formation of runaway massive objects has previously been suggested through an analysis of observational properties of both NSCs and Massive Black Holes by Escala (2021), from the comparison of collision timescales with ages in galactic nuclei. The scenario was further tested with satisfactory results, using a set of idealized N-body simulations employing constant stellar masses by Vergara et al. (2023), where a critical mass for collisions was introduced for the comparison of the collision time with the age of the system.
@@
Here we have generalized this concept to take into account the presence of an external potential, and we employed the result from a large range of numerical simulations and observational data from the literature. This way we compared with the previous results and determined how the efficiency for the formation of massive objects depends on the ratio of initial stellar mass over the critical mass scale defined in Eq. 2. This included simulations with different types of radial profiles in their initial conditions (e.g. Plummer, 1911; King, 1966; Miyamoto & Nagai, 1975), simulations employing different stellar IMFs (e.g. Salpeter, 1955; Scalo, 1986; Kroupa, 2001) as well as simulations including stellar evolution (Hurley et al., 2000, 2002), including simulations with one million stars (Panamarev et al., 2019; Arca Sedda et al., 2023a, b, c).
@@
The comparison of these data is not completely trivial, for example, not all the output is always reported in the literature, and while the initial conditions are usually well-described, the final stellar mass in the cluster and the mass loss due to escapers are not always given, but forms an important ingredient in our efficiency parameter $\epsilon_{\mathrm{BH}}$. Nonetheless, even under the simplifying assumption that mass loss would be negligible, our analysis has shown a clear correlation between the efficiency to form a massive object and the ratio of initial stellar mass over critical mass. Particularly, when that ratio is larger than 1, the typical efficiencies lie at least in the range of $20-50\%$, while for lower ratios the most frequent outcome is within the percent range. We note that there is scatter in this relation and we also note some differences in the different sets of simulations, which also is unsurprising given the different assumptions entering them. We further note that the obtained correlation improves if we derive a correction factor from the simulations of Reinoso et al. (2020); Vergara et al. (2021, 2023), as for those simulations we know the fraction of stellar mass lost due to escapers. Assuming that the physics are similar in other simulations, the correlation becomes more clear and we find efficiencies of $20-60\%$ when the ratio of initial stellar over critical mass is larger than $1$.
@@
Subsequently, we have analyzed observational data of dense stellar systems, including GCs, NSCs, and UCDs. The GCs are perhaps the clearest case as these are very old systems without any recent star formation and where the general evolution is very well understood. For these cases, the ratio of initial stellar over critical mass is in the range of $0.005-0.03$, and the efficiencies are very low $\sim 5\%$. The simulations of Portegies Zwart et al. (1999); Fujii & Portegies Zwart (2013); Mapelli (2016) and Arca Sedda et al. (2023a, b, c) show low $\epsilon_{\rm BH}$ and further substantiated by observation of Harris (e.g 1996) and Lützgendorf et al. (e.g 2011, 2013a, 2013b, 2013c). This supports the scenario of IMBHs formation in GCs.
@@
In the case of NSCs, we notice more scatter, as the systems are more complex including complicated star formation histories and typically also mergers with other stellar systems. They cover a wide range of initial over critical mass ratios of $0.01-100$, and despite the presence of scatter, it is visible that very high efficiencies above $50\%$ are only reached for mass ratios above $0.3$.
@@
The UCDs are perhaps the most uncertain systems, as their nature and evolution are not fully clear. We have explored assumptions of these systems both assuming an evolution more like a NSCs and alternatively using an evolutionary path with a fixed radius but including an external potential due to dark matter. The first case is more favorable while the second somewhat increases the scatter, though both cases do not fundamentally change the picture. It is however noticeable that the observational data do include more scatter than the simulations. This could be due to the higher complexity of those systems in general, and of course, it is also not guaranteed that the observed systems formed just as a result of collisions, but it could be that additional gas physics was also involved, which could help to increase the efficiency to form a massive object (e.g. Schleicher et al., 2022).
@@
Overall, while it is clear that additional factors such as the complex history in real stellar systems, and the possible presence of gas, most likely have an influence on the results and may also be able to explain the scatter that we observe, we find efficiencies of $30-100\%$, which are correlated with ratios of initial stellar over critical mass above 1. Observations reach higher efficiencies due to the secular evolutionary process of evaporation of the cluster. This result should not be particularly surprising, as it only means that the collision timescale should be short enough to play a relevant role within a given stellar system, but should be useful to determine the potential of a system to form very massive objects. It is worth stressing out that high efficiencies can be reached for a wide range of systems. This opens the potential to form quite massive black holes via the collision channel in massive NSCs as long as their stellar mass is larger than their critical mass.
@@
In the future, it will be useful for this scenario to be tested further. For example, at the moment only a few N-body simulations with one million stars are available, such as Panamarev et al. (2019); Arca Sedda et al. (2023a, b, c). More simulations and more compact systems will allow us to test this scenario further and under more realistic conditions. At the same time, the complexity of simulations could still be increased, by implementing ongoing star formation as expected within real NSCs, and/or mergers with GCs as another possible growth channel. To understand the importance of this black hole formation mechanism, statistical predictions will be important in combining semi-analytic galaxy evolution models as Sassano et al. (e.g. 2021); Trinca et al. (e.g. 2022) with black hole formation recipes employing the concept of the critical mass scale to obtain statistical predictions both at high and low redshift. In the local Universe, such predictions could be compared to the black hole mass function determined via the James Webb Space Telescope (JWST) [^2] Webpage JWST: https://webb.nasa.gov/ and the Extremely Large Telescope (ELT) [^3] Webpage ELT: https://elt.eso.org/, where new IMBHs may be discovered at the lower mass end and in general more data will become available for testing the scenario proposed here. But at high redshift, this scenario can operate, and its relevance for the formation of high redshift black holes needs to be further explored.
@@
5.While the potential importance of the collision-based channels for black hole formation was already suggested by Rees (1984), previous estimates typically suggest they form black hole masses up to the order of $10^{3}$ M (e.g. Devecchi & Volonteri, 2009; Devecchi et al., 2010; Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018). The results obtained here considerably strengthen and support the relevance of the collision-based channel, suggesting that it could efficiently operate in dense NSCs and the efficiency of SMBH formation in typical NSCs appears roughly consistent with this scenario, including SMBH masses of order $10^{6}$ M and somewhat above. It is even further consistent with the very low efficiencies in GCs, while the results for UCDs are not fully conclusive, as they depend on some assumptions to be made concerning these objects. What we can conclude from the comparison with the large set of simulations analyzed in this work is that the scenario holds over a wide range of possible conditions, including variations of the radial profile of the clusters, the presence of an initial mass function as well as the inclusion of additional physics such as stellar evolution. Future simulations are going to include high particle numbers and exploring additional physics, including improved treatments of collisions and stellar physics (e.g. Kamlah et al., 2024), but also investigating the role of rotation and binaries will allow us to probe this scenario in even more detail.
@@ Acknowledgements.
We thank the anonymous referee for his/her comments, which helped to improve the manuscript quality. MCV acknowledges funding through ANID (Doctorado acuerdo bilateral DAAD/62210038) and DAAD (funding program number 57600326). DRGS, AE and ML acknowledge the center for Astrophysics and Associated Technologies CATA (FB210003) and via ANID Fondo 2022 QUIMAL 220002. DRGS and ML also acknowledge financial support from FONDECYT Regular grant ##\##1201280 and through the Alexander von Humboldt - Foundation, Bonn, Germany. BR acknowledges funding through ANID (CONICYT-PFCHA/Doctorado acuerdo bilateral DAAD/62180013), DAAD (funding program number 57451854), and the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). FFD, AWHK, and RS acknowledge financial support by German Science Foundation (DFG) project Sp 345/22-1 and Sp 345/24-1. Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility, and we also acknowledge the Gauss center for Supercomputing e.V. for computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Booster at Jülich Supercomputing center (JSC). These resources made the presented work possible, by supporting its development. NH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). NH acknowledges funding through https://www.lacegal.com/LACEGAL, a Latinamerican Chinese European Galaxy Formation Network. This project has received funding from the European Union’s HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Sklodowska-Curie grant agreement number 101086388.
@@ Appendix A Computation of the critical mass and black hole formation efficiency for simulations
Fujii & Portegies Zwart (2013) delve into the role of stellar collisions in ensemble star clusters and virialized solo-cluster models, in this work, we focus on the last ones. They use a Salpeter (1955) IMF with $M_{\rm min}=1\mathrm{M_{\odot}}$ and $M_{\rm max}=100\mathrm{M_{\odot}}$ and density profile of King (1966). To compute the critical mass and the black hole formation efficiency we use the number of stars ($N$), the cluster mass ($M_{\mathrm{i}}$), the half mass radius, the mean stellar mass ($\overline{m_{*}}=M_{\mathrm{i}}/N$), and finally the black hole mass after a simulation time of 555\,5Myr. In the study of Sakurai et al. (2017) they explore runaway collisions in the first star cluster, using a Salpeter (1955) IMF varying the masses between $3-100\mathrm{M_{\odot}}$. The distribution of stars does not follow any profile mentioned before, since they obtain the positions of the stars from previous hydrodynamics simulations. The parameters adopted to compute the $M_{\mathrm{crit}}$ and $\epsilon_{\mathrm{BH}}$ were the simulation time ($\tau=3\,$Myr), the initial cluster, the number of stars and the black hole mass.
Portegies Zwart et al. (1999) explored runaway collisions in young compact star clusters. They employed a Scalo (1986) IMF, which varies the stellar mass between $0.1$ and $100\,M_{\odot}$. We use $N$, $\overline{m_{*}}$, $M_{\mathrm{i}}=N*\overline{m_{*}}$, and $R_{\mathrm{h}}$ besides the values for the black hole mass and the simulation time were extracted from Figure 2 to compute $M_{\text{crit}}$ and $\epsilon_{\text{BH}}$. Mapelli (2016) explore the impact of metallicity in runaway collisions in young dense star clusters. The authors use a Kroupa (2001) IMF with $M_{\rm min}=0.1\mathrm{M_{\odot}}$ and $M_{\rm max}=150\mathrm{M_{\odot}}$. The parameters used to calculate the critical mass and the black hole formation efficiency are $N$, $M_{\mathrm{i}}$, $\overline{m_{*}}=M_{\mathrm{i}}/N$, and $R_{\mathrm{h}}$. The stellar system was evolved for over 171717\,17Myr, forming binary systems. We just consider the mass of the most massive one of the binary system. Both works use a King (1966) profile.
@@
Panamarev et al. (2019) simulated the galactic center of the Milky Way, using one million stars, including a Kroupa (2001) IMF with masses ranging from $0.08$ to $100\,M_{\odot}$. To assess the efficiency of black hole formation and the critical mass, we used an initial mass, the half-mass radius, the time ($\tau=5.5\,\text{Gyr}$), and the black hole mass was taken from their Figure 7. In a different study, Vergara et al. (2023) investigated runaway collisions in NSCs, employing models with equal-mass stars distributed according to a Plummer distribution (Plummer 1911). To determine the critical mass and black hole formation efficiency, we considered the initial cluster mass, virial radii, the stellar mass and radius, and the mass of very massive stars formed after $\tau=10\,\text{Myr}$, as Reinoso et al. (2018, 2020); Vergara et al. (2021), they assume that very massive stars evolve enough time to form a black hole. We computed the half-mass radius using the same assumptions as Reinoso et al. (2020) and Katz et al. (2015). Besides Arca Sedda et al. (2023a, b, c) simulated star cluster using $(0.12,0.3,0.6,1)\times 10^{6}$ stars, including a Kroupa (2001) IMF ranging from $0.08$ to $150\,M_{\odot}$. These simulations also include primordial binaries up to $33\%$ of the total number of stars. The simulations last around a few gigayears. We took the maximum black hole mass formed in the stellar system (see their Table 1).
Rizzuto et al. (2023) explore tidal disruption/capture runaway collisions in dense clusters, including post-Newtonian effects. The use of a Kroupa (2001) IMF, with a range between $0.08$ to $2\,M_{\odot}$, the low-mass stars were chosen to avoid stellar evolution. They also include an initial black hole within the cluster with initial masses of $50$, $300$ and, $2000\,M_{\odot}$. Almost all the simulations run for times around $40$  Myr.
To compute the efficiency of black hole formation and the critical mass, we used an initial mass, the half-mass radius, the simulation time, and the black hole mass.

### v2

@@ 4 Summary and conclusions
In this paper, we investigate how the efficiency of black hole formation via collisions in stellar systems depends on the ratio of stellar over critical mass. The importance of the collision timescale in the formation of runaway massive objects has previously been suggested through an analysis of observational properties of both NSCs and massive black holes by Escala (2021), from the comparison of collision timescales with ages in galactic nuclei. The scenario was further tested with satisfactory results, using a set of idealized N-body simulations employing constant stellar masses by Vergara et al. (2023), where a critical mass for collisions was introduced for the comparison of the collision time with the age of the system.
@@
Here, we have generalized this concept to take into account the presence of an external potential and we employed the result from a large range of numerical simulations and observational data from the literature. This way we compared with the previous results and determined how the efficiency for the formation of massive objects depends on the ratio of initial stellar mass over the critical mass scale defined in Eq. 2. This included simulations with different types of radial profiles in their initial conditions (e.g., Plummer, 1911; King, 1966; Miyamoto & Nagai, 1975), simulations employing different stellar IMFs (e.g., Salpeter, 1955; Scalo, 1986; Kroupa, 2001) as well as simulations including stellar evolution (Hurley et al., 2000, 2002) and simulations with one million stars (Panamarev et al., 2019; Arca Sedda et al., 2023a, b, c).
@@
The comparison of these data is not completely trivial, for example, not all the output is always reported in the literature; also, while the initial conditions are usually well-described, the final stellar mass in the cluster and the mass loss due to escapers are not always given, but this still forms an important ingredient in our efficiency parameter, $\epsilon_{\mathrm{BH}}$. Nonetheless, even under the simplifying assumption that mass loss would be negligible, our analysis has shown a clear correlation between the efficiency required to form a massive object and the ratio of initial stellar mass over critical mass. Particularly, when that ratio is larger than 1, the typical efficiencies lie at least in the range of $20-50\%$; whereas for lower ratios, the most frequent outcome is within the percent range. We note that there is scatter in this relation and we also note some differences in the different sets of simulations, which are also unsurprising given the different assumptions involved. We further note that the obtained correlation improves if we derive a correction factor from the simulations of Reinoso et al. (2020); Vergara et al. (2021, 2023), as for those simulations, we know the fraction of stellar mass lost due to escapers. Assuming that the physics are similar in other simulations, the correlation becomes more clear and we find efficiencies of $20-60\%$ when the ratio of initial stellar over critical mass is greater than $1$.
@@
Subsequently, we have analyzed observational data of dense stellar systems, including GCs, NSCs, and UCDs. The GCs are perhaps the clearest case as these are very old systems without any recent star formation and where the general evolution is very well understood. For these cases, the ratio of initial stellar over critical mass is in the range of $0.005-0.03$ and the efficiencies are very low $\sim 5\%$. The simulations of Portegies Zwart et al. (1999); Fujii & Portegies Zwart (2013); Mapelli (2016) and Arca Sedda et al. (2023a, b, c) show low $\epsilon_{\rm BH}$ and further substantiated by observation of Harris (e.g 1996) and Lützgendorf et al. (e.g 2011, 2013a, 2013b, 2013c). This supports the scenario of IMBHs formation in GCs.
@@
In the case of NSCs, we notice more scatter, as the systems are more complex including complicated star formation histories and typically also mergers with other stellar systems. They cover a wide range of initial over critical mass ratios of $0.01-100$ and, despite the presence of scatter, it is visible that very high efficiencies above $50\%$ are only reached for mass ratios above $0.3$.
@@
The UCDs are perhaps the most uncertain systems, as their nature and evolution are not fully clear. We have explored assumptions of these systems both assuming an evolution more like a NSCs and alternatively using an evolutionary path with a fixed radius but including an external potential due to dark matter. The first case is more favorable, while the second somewhat increases the scatter, although neither of these cases fundamentally change the picture. However, it is noticeable that the observational data do include more scatter than the simulations. This could be due to the higher complexity of those systems in general. Also, it is certainly not guaranteed that the observed systems formed just as a result of collisions, but it could be the case that additional gas physics was also involved, which could help to increase the efficiency with respect to forming a massive object (e.g., Schleicher et al., 2022).
@@
Overall, while it is clear that additional factors such as the complex history in real stellar systems and the possible presence of gas most likely have an influence on the results and may also be able to explain the scatter that we observe, we find efficiencies of $30-100\%$; these results are correlated with ratios of initial stellar over critical mass above 1. Observations reach higher efficiencies due to the secular evolutionary process of evaporation of the cluster. This result should not be particularly surprising, as it only means that the collision timescale should be short enough to play a relevant role within a given stellar system, but should be useful to determine the potential of a system to form very massive objects. It is worth stressing that high efficiencies can be reached for a wide range of systems. This opens up the potential to form quite massive black holes via the collision channel in massive NSCs, as long as their stellar mass is larger than their critical mass.
@@
In the future, it will be useful for this scenario to be tested further. For example, there are only a few N-body simulations with one million stars available, such as those of Panamarev et al. (2019); Arca Sedda et al. (2023a, b, c). More simulations and more compact systems will allow us to test this scenario further and under more realistic conditions. At the same time, the complexity of simulations could still be increased, by implementing ongoing star formation as expected within real NSCs and/or mergers with GCs as another possible growth channel. To understand the importance of this black hole formation mechanism, statistical predictions will be important in combining semi-analytic galaxy evolution models, as those of Sassano et al. (e.g. 2021); Trinca et al. (e.g. 2022), with black hole formation recipes employing the concept of the critical mass scale to obtain statistical predictions at both high and low redshift. In the Local Universe, such predictions could be compared to the black hole mass function determined via the James Webb Space Telescope (JWST) and the Extremely Large Telescope (ELT), where new IMBHs may be discovered at the lower mass end and in general more data will become available for testing the scenario proposed here. At high redshift, this scenario can operate and its relevance for the formation of high redshift black holes needs to be further explored.
@@
While the potential importance of the collision-based channels for black hole formation has already been suggested by Rees (1984), previous estimates typically insinuate that they form black hole masses up to the order of $10^{3}$ M (e.g., Devecchi & Volonteri, 2009; Devecchi et al., 2010; Katz et al., 2015; Sakurai et al., 2017; Reinoso et al., 2018). The results obtained here considerably strengthen and support the relevance of the collision-based channel, suggesting that it could efficiently operate in dense NSCs and the efficiency of SMBH formation in typical NSCs appears roughly consistent with this scenario, including SMBH masses on the order of $10^{6}$ M and somewhat above. It is even further consistent with the very low efficiencies in GCs, while the results for UCDs are not fully conclusive, as they depend on certain assumptions being made concerning these objects. What we can conclude from the comparison with the large set of simulations analyzed in this work is that the scenario holds over a wide range of possible conditions, including variations of the radial profile of the clusters, the presence of an initial mass function, and the inclusion of additional physics such as stellar evolution. Future simulations with high particle numbers and exploring additional physics, including improved treatments of collisions and stellar physics (e.g., Kamlah et al., 2024), and investigating the role of rotation and binaries will allow us to probe this scenario in even greater detail.
@@ Acknowledgements.
We thank the anonymous referee for his/her comments, which helped to improve the manuscript quality. MCV acknowledges funding through ANID (Doctorado acuerdo bilateral DAAD/62210038) and DAAD (funding program number 57600326). DRGS, AE and ML acknowledge the center for Astrophysics and Associated Technologies CATA (FB210003) and via ANID Fondo 2022 QUIMAL 220002. DRGS and ML also acknowledge financial support from FONDECYT Regular grant ##\##1201280 and through the Alexander von Humboldt - Foundation, Bonn, Germany. BR acknowledges funding through ANID (CONICYT-PFCHA/Doctorado acuerdo bilateral DAAD/62180013), DAAD (funding program number 57451854), and the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). FFD, AWHK, and RS acknowledge financial support by German Science Foundation (DFG) project Sp 345/22-1 and Sp 345/24-1. Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility, and we also acknowledge the Gauss center for Supercomputing e.V. for computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Booster at Jülich Supercomputing center (JSC). These resources made the presented work possible, by supporting its development. NH is a fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). NH acknowledges funding through LACEGAL, a Latinamerican Chinese European Galaxy Formation Network. This project has received funding from the European Union’s HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Sklodowska-Curie grant agreement number 101086388.
@@ Appendix A Computation of the critical mass and black hole formation efficiency for simulations
Fujii & Portegies Zwart (2013) delved into the role of stellar collisions in ensemble star clusters and virialized solo-cluster models. In this work, we focus on the latter. Those authors used a Salpeter (1955) IMF with $M_{\rm min}=1\mathrm{M_{\odot}}$ and $M_{\rm max}=100\mathrm{M_{\odot}}$ and density profile of King (1966). To compute the critical mass and the black hole formation efficiency, we used the number of stars ($N$), the cluster mass ($M_{\mathrm{i}}$), the half-mass radius, the mean stellar mass ($\overline{m_{*}}=M_{\mathrm{i}}/N$), and, finally, the black hole mass after a simulation time of 555\,5Myr. In the study of Sakurai et al. (2017) they explore runaway collisions in the first star cluster, using a Salpeter (1955) IMF varying the masses between $3-100\mathrm{M_{\odot}}$. The distribution of stars does not follow any profile mentioned before, since the positions of the stars were obtained from previous hydrodynamics simulations. The parameters adopted to compute the $M_{\mathrm{crit}}$ and $\epsilon_{\mathrm{BH}}$ were the simulation time ($\tau=3\,$Myr), the initial cluster, the number of stars, and the black hole mass.
Portegies Zwart et al. (1999) explored runaway collisions in young compact star clusters. They employed a Scalo (1986) IMF, which varies the stellar mass between $0.1$ and $100\,M_{\odot}$. We used $N$, $\overline{m_{*}}$, $M_{\mathrm{i}}=N*\overline{m_{*}}$, and $R_{\mathrm{h}}$ in addition to the values for the black hole mass and the simulation time (from Figure 2) to compute $M_{\text{crit}}$ and $\epsilon_{\text{BH}}$. Mapelli (2016) explored the impact of metallicity in runaway collisions in young dense star clusters. The authors used the Kroupa (2001) IMF with $M_{\rm min}=0.1\mathrm{M_{\odot}}$ and $M_{\rm max}=150\mathrm{M_{\odot}}$. The parameters used to calculate the critical mass and the black hole formation efficiency are $N$, $M_{\mathrm{i}}$, $\overline{m_{*}}=M_{\mathrm{i}}/N$, and $R_{\mathrm{h}}$. The stellar system was evolved for over 171717\,17Myr, forming binary systems. We just consider the mass of the most massive one of the binary system. Both works used the King (1966) profile.
@@
Panamarev et al. (2019) simulated the galactic center of the Milky Way, using one million stars, including a Kroupa (2001) IMF with masses ranging from $0.08$ to $100\,M_{\odot}$. To assess the efficiency of black hole formation and the critical mass, we used an initial mass, half-mass radius, time ($\tau=5.5\,\text{Gyr}$), and the black hole mass taken from their Figure 7. In a different study, Vergara et al. (2023) investigated runaway collisions in NSCs, employing models with equal-mass stars distributed according to a Plummer distribution (Plummer 1911). To determine the critical mass and black hole formation efficiency, we considered the initial cluster mass, virial radii, stellar mass and radius, and the mass of very massive stars formed after $\tau=10\,\text{Myr}$. In Reinoso et al. (2018, 2020); Vergara et al. (2021), the authors assumed that very massive stars evolve enough time to form a black hole. We computed the half-mass radius using the same assumptions as Reinoso et al. (2020) and Katz et al. (2015). Besides Arca Sedda et al. (2023a, b, c) simulated star cluster using $(0.12,0.3,0.6,1)\times 10^{6}$ stars, including a Kroupa (2001) IMF ranging from $0.08$ to $150\,M_{\odot}$. These simulations also include primordial binaries up to $33\%$ of the total number of stars. The simulations last around a few gigayears. We took the maximum black hole mass formed in the stellar system (see their Table 1).
Rizzuto et al. (2023) explored tidal disruption and capture runaway collisions in dense clusters, including post-Newtonian effects. The use of a Kroupa (2001) IMF, with a range between $0.08$ to $2\,M_{\odot}$, the low-mass stars were chosen to avoid stellar evolution. They also include an initial black hole within the cluster with initial masses of $50$, $300,$ and, $2000\,M_{\odot}$. Almost all the simulations run for times around $40$  Myr. To compute the efficiency of black hole formation and the critical mass, we used an initial mass, the half-mass radius, the simulation time, and the black hole mass.